OpenStudy (anonymous):

WILL GIVE MEDAL FOR RIGHT ANSWER: A rectangular storage container with an open top is to have a volume of 26 cubic meters. The length of its base is twice the width. Material for the base costs 12 dollars per square meter. Material for the sides costs 9 dollars per square meter. Find the cost of materials for the cheapest such container.

OpenStudy (camper4834):

i see L=2W B=2W*W

OpenStudy (camper4834):

V=2W^2 * H

OpenStudy (camper4834):

26=2W^2 * H Cost = 54HW+24W^2

OpenStudy (camper4834):

Cost = 702/w + 24w^2

OpenStudy (camper4834):

derivative = 0

OpenStudy (camper4834):

\[702w^{-1}+24w^2\] derivative = \[-702w^{-2}+48w\]

OpenStudy (camper4834):

\[-\frac{ 702 }{ w^{2} }+48w=0\]\[48w=\frac{ 702 }{ w^{2} }\]\[w=\sqrt[3]{\frac{ 117 }{ 8 }}\]

OpenStudy (camper4834):

\[w \approx 2.4454866232544\]

OpenStudy (camper4834):

the least amount of cost is when L = 4.8909732465088 W= 2.4454866232544 H =2.1737658873373